OpenStudy (anonymous):

A missle is launched from the ground. Its height h(x), can be represented by a quadratic function in terms of time, x, in seconds. After 1 second, the missle is 130 feet in the air; after 2 seconds it is 240 feet in the air. Find the height, in feet, of the missle after 11 seconds in the air.

OpenStudy (whpalmer4):

Personally, I think t is a better letter for time than x :-) \[h(t) = at^2 + bt + c\] \[h(0) = 0\]\[h(1) = 130\]\[h(2) = 240\] So now write the 3 equations you can form from this information and solve for the values of a, b, and c.

OpenStudy (anonymous):

@whpalmer4 Thats what I dont understand... how to plug then into an equation.

OpenStudy (whpalmer4):

at t=0: \[h(t) = 0 = a(0)^2 + b(0) + c \rightarrow c = 0 \] at t=1: \[h(t) = 130 = a(1)^2 + b(1) + 0 = a + b\] at t=2: \[h(t) = 240 = a(2)^2 + b(2) + 0 = 4a + 2b\] That gives us 2 equations in 2 unknowns: \[a + b = 130\]\[4a+2b = 240\] Can you solve that?

OpenStudy (anonymous):

Got it! I'll solve it. thanks

OpenStudy (whpalmer4):

Let me know what you get.

OpenStudy (whpalmer4):

I had to do it twice, because I accidentally wrote 260 instead of 240 :-)